1. For every person who walked there were 44 who drove. If 270 made the trip, how many walked?
2. Jenna pedaled the 24 miles to the lake against the wind and averaged 8 miles per hour on the return trip she averaged 12 miles per hour. Find Jenna's average speed for the round trip.

Answer by stanbon(75887)

1. For every person who walked there were 44 who drove. If 270 made the trip, how many walked?
-----
Equation:
x + 44x = 270
45x = 270
x = 6 (# who walked)
---------------------------
2. Jenna pedaled the 24 miles to the lake against the wind and averaged 8 miles per hour on the return trip she averaged 12 miles per hour. Find Jenna's average speed for the round trip.
-----
To Lake DATA:
dist = 24 miles ; rate = 8 mph ; time = 24/8 = 3 hrs
-------------
From Lake DATA:
dist = 24 miles ; rate = 12 mph ; time = 24/12 = 2 hrs
------
Average speed = total distance/total time = 48 miles/5 hrs = 9.6 mph
--------------------
Cheers,
Stan H.
---------------------